OpenStudy (anonymous):

solve 2x^2+3x=-5 solve 2x^2+3x=-5 @Mathematics

OpenStudy (mimi_x3):

2x^2 + 3x + 5 = 0 a = 2 , b = 3 , c = 5 Sub it into the quadratic formula

myininaya (myininaya):

\[x^2+\frac{3}{2}x=\frac{-5}{2}\] \[x^2+\frac{3}{2}x+(\frac{3}{2 \cdot 2})^2=\frac{-5}{2}+(\frac{3}{2 \cdot 2})^2\] \[x^2+\frac{3}{2}x+(\frac{3}{4})^2=\frac{-5}{2}+\frac{9}{16}\] \[(x+\frac{3}{4})^2=\frac{-40+9}{16}\]

myininaya (myininaya):

\[x+\frac{3}{4}= \pm \sqrt{\frac{-31}{16}}\]

myininaya (myininaya):

\[x=-\frac{3}{4} \pm \frac{\sqrt{-31}}{4}=\frac{-3 \pm i \sqrt{31}}{4}\]

OpenStudy (agreene):

2x^2+3x=-5 x^2+3x/2 = -5/2 add 9/16 to both sides so we can factor x^2+3x/2+9/16=-31/16 (x+3/4)^2=-31/16 \[|x+\frac{3}{4}|=\sqrt{-\frac{31}{16}}\] \[|x+\frac{3}{4}|=\frac{i \sqrt{31}}{4}\] \[x+\frac{3}{4}=\pm\frac{i \sqrt{31}}{4}\] \[x=-\frac{1}{4}i(\sqrt{31}-3i)\] and \[x=\frac{1}{4}i(\sqrt{31}+3i)\]